# ADBM: Adversarial Diffusion Bridge Model for Reliable Adversarial Purification

### v1

@@ B.2 The Proof of Theorem 1
Given an adversarial example ${\mathbf{x}}_{0}^{a}$ and assuming the training loss $L_{b}\leq\delta$, the distance between the purified example of ADBM and the clean example ${\mathbf{x}}_{0}$, denoted as $\|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\|$, is bounded by $\delta$ in expectation (constant omitted) when using a one-step DDIM sampler. Specifically, we have $\mathbb{E}_{{\boldsymbol{\epsilon}}}[\|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\|^{2}]\leq\frac{(1-\bar{\alpha}_{T})T}{{\bar{\alpha}_{T}}}\delta$, where $\frac{(1-\bar{\alpha}_{T})T}{{\bar{\alpha}_{T}}}$ is the constant.
@@
This inequality holds when we use the DDIM reverse sampler and set the reverse step $s=1$. According to Song et al. [22], the reverse process of DDIM is
$\hat{{\mathbf{x}}}_{\tau_{i-1}}=\sqrt{\bar{\alpha}_{\tau_{i-1}}}\left(\frac{\hat{{\mathbf{x}}}_{\tau_{i}}-\sqrt{1-\bar{\alpha}_{\tau_{i}}}{\boldsymbol{\epsilon}}_{\theta}(\hat{{\mathbf{x}}}_{\tau_{i}},\tau_{i})}{\sqrt{\bar{\alpha}_{\tau_{i}}}}\right)+\sqrt{1-\bar{\alpha}_{\tau_{i-1}}}{\boldsymbol{\epsilon}}_{\theta}(\hat{{\mathbf{x}}}_{\tau_{i}},\tau_{i}),$ (23)
where $\{\tau_{0},\ldots,\tau_{s}\}$ is a linearly increasing sub-sequence of $[0,\ldots,t]$, $\tau_{0}=1,\tau_{s}=t$. Setting $s=1$, then it yields:
$\hat{{\mathbf{x}}}_{0}=\sqrt{\bar{\alpha}_{0}}\left(\frac{{\mathbf{x}}_{t}^{a}-\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{a},t)}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{0}}{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{a},t).$ (24)
Since $\sqrt{\bar{\alpha}_{0}}=1$, ${\mathbf{x}}_{t}^{a}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}$, ${\mathbf{x}}_{0}^{a}={\mathbf{x}}_{0}+{\boldsymbol{\epsilon}}_{a}$, and $t=T$, Eq. 24 can be written as
$\hat{{\mathbf{x}}}_{0}$ $={\mathbf{x}}_{0}^{a}+\frac{\sqrt{1-\bar{\alpha}_{T}}}{\sqrt{\bar{\alpha}_{T}}}({\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{a},T))$ (25)
$={\mathbf{x}}_{0}+\frac{\sqrt{1-\bar{\alpha}_{T}}}{\sqrt{\bar{\alpha}_{T}}}(\frac{\sqrt{\bar{\alpha}_{T}}}{\sqrt{1-\bar{\alpha}_{T}}}{\boldsymbol{\epsilon}}_{a}+{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{a},T)).$
@@
$\|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\|$ $=\frac{\sqrt{1-\bar{\alpha}_{T}}}{\sqrt{\bar{\alpha}_{T}}}\left\|\frac{\sqrt{\bar{\alpha}_{T}}}{\sqrt{1-\bar{\alpha}_{T}}}{\boldsymbol{\epsilon}}_{a}+{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{a},T)\right\|$ (26)
$=\frac{\sqrt{1-\bar{\alpha}_{T}}}{\sqrt{\bar{\alpha}_{T}}}\left\|\frac{\sqrt{\bar{\alpha}_{T}}}{\sqrt{1-\bar{\alpha}_{T}}}{\boldsymbol{\epsilon}}_{a}+{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{d},T)\right\|,$
where the second equivalence holds due to $t=T$ and ${\mathbf{x}}_{T}^{d}={\mathbf{x}}_{T}^{a}-k_{T}{\boldsymbol{\epsilon}}_{a}={\mathbf{x}}_{T}^{a}$. Considering that
$L_{b}$ $=\mathbb{E}_{{\boldsymbol{\epsilon}},t}\left[\left\|\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\cdot\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{a}+{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t)\right\|^{2}\right]$ (27)
$=\frac{1}{T}\sum_{t=0}^{T}\mathbb{E}_{{\boldsymbol{\epsilon}}}\left[\left\|\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\cdot\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{a}+{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t)\right\|^{2}\right]$
@@
$\mathbb{E}_{{\boldsymbol{\epsilon}}}\left[\frac{{1-\bar{\alpha}_{T}}}{{\bar{\alpha}_{T}}}\left\|\frac{\sqrt{\bar{\alpha}_{T}}}{\sqrt{1-\bar{\alpha}_{T}}}{\boldsymbol{\epsilon}}_{a}+{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{d},T)\right\|^{2}\right]\leq T\cdot\delta$ (28)
@@
$\mathbb{E}_{{\boldsymbol{\epsilon}}}\left[\left|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\right\|^{2}\right]$ $=\mathbb{E}_{{\boldsymbol{\epsilon}}}\left[\frac{{1-\bar{\alpha}_{T}}}{{\bar{\alpha}_{T}}}\left\|\frac{\sqrt{\bar{\alpha}_{T}}}{\sqrt{1-\bar{\alpha}_{T}}}{\boldsymbol{\epsilon}}_{a}+{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{d},T)\right\|^{2}\right]$ (29)
@@ B.3 The Proof of Theorem 2
The concept of infinite timestep can be viewed as dividing a finite length of time into infinitesimal intervals, which corresponds to the situation of the following SDEs proposed by Song et al. [23]. Denoting ${\mathbf{w}}$ as the standard Wiener process, $\bar{{\mathbf{w}}}$ as the reverse-time standard Wiener process, and $p_{t}({\mathbf{x}})$ the probability density of ${\mathbf{x}}_{t}$, the forward process can be described by
@@
$\propto\int\mathbbm{1}_{\{{\mathbf{x}}_{0}\notin{\mathbb{D}}_{a}\}}\exp\left(-\frac{\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\boldsymbol{\epsilon}}_{a}\|^{2}}{2(1-\bar{\alpha}_{t})}\right)p({\mathbf{x}}_{0}){\mathrm{d}}{\mathbf{x}}_{0}$
@@
$\left[\exp\left(-\frac{\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\boldsymbol{\epsilon}}_{a}\|_{2}^{2}}{2(1-\bar{\alpha}_{t})}\right)-\exp\left(-\frac{\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}\|_{2}^{2}}{2(1-\bar{\alpha}_{t})}\right)\right]{\mathrm{d}}{\mathbf{x}}_{0}.$
Therefore, we only need to compare $\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\boldsymbol{\epsilon}}_{a}\|_{2}^{2}$ and $\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}\|_{2}^{2}$ for comparing $P(B)$ and $P(D)$.
Since ${\mathbf{x}}_{t}^{a}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}$ and ${\mathbf{x}}_{0}^{a}={\mathbf{x}}_{0}+{\boldsymbol{\epsilon}}_{a}$,
$\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\boldsymbol{\epsilon}}_{a}\|$ $=\|\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\boldsymbol{\epsilon}}_{a}\|$ (38)
$=\|\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}+k_{t}{\boldsymbol{\epsilon}}_{a}\|,$
$\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}\|$ $=\|\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}\|$ (39)
$=\|\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}+\sqrt{\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{a}\|.$
@@
$\mathbb{E}_{{\boldsymbol{\epsilon}}}\|\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}+k_{t}{\boldsymbol{\epsilon}}_{a}\|<\mathbb{E}_{{\boldsymbol{\epsilon}}}\|\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}+\sqrt{\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{a}\|.$ (40)
@@
$\mathbb{E}_{{\boldsymbol{\epsilon}}}\left[\exp\left(-\frac{\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\boldsymbol{\epsilon}}_{a}\|_{2}^{2}}{2(1-\bar{\alpha}_{t})}\right)\right]>\mathbb{E}_{{\boldsymbol{\epsilon}}}\left[\exp\left(-\frac{\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}\|_{2}^{2}}{2(1-\bar{\alpha}_{t})}\right)\right],$ (41)
@@
Appendix C Additional Settings
C.1 Additional Training Settings of ADBM
We implemented ADBM based loosely on the original implementations by Song et al. [23]. Note that when fine-tuning the ADBM with the pre-trained checkpoint, following Song et al. [23], we also adopted the continuous version of Eq. 9. This version is conceptually similar to its discrete counterpart, with the exception that $t$ in Eq. 9 represents a continuous value rather than a discrete one. Moreover, we used the Adam optimizer [53] and incorporated the exponential moving average of models, with the average rate being 0.999. The batch size was set to 128 for SVHN and CIFAR-10, 112 for Tiny-ImageNet, and 64 for ImageNet-100 (due to memory constraints). All experiments were run using PyTorch 1.12.1 and CUDA 11.3 on 4 NVIDIA 3090 GPUs.
C.2 Additional Adaptive Attack Configurations
When performing the adaptive attacks (PGD with EOT) for the AP methods, we set the step sizes to be 0.007, 0.5, and 0.005 for $l_{\infty}$, $l_{1}$, and $l_{2}$ attacks, respectively. For the $l_{1}$ attack, we set the sparsity level as 0.95. As shown in Sec. 3, the adaptive evaluating is quite time-consuming. Thus following Nie et al. [16], we conducted the adaptive attacks three times on a subset of 512 randomly sampled images from the test dataset and reported the mean accuracy along with the standard deviation. We note that sometimes the standard deviation is omitted to present the results more clearly.

### v3

@@ C.2 The Proof of Theorem 1
Given an adversarial example ${\mathbf{x}}_{0}^{a}$ and assuming the training loss $L_{b}\leq\delta$, the distance between the purified example of ADBM and the clean example ${\mathbf{x}}_{0}$, denoted as $\|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\|$, is bounded by $\delta$ in expectation (constant omitted) when using a one-step DDIM sampler. Specifically, we have $\mathbb{E}_{{\bm{\epsilon}}}[\|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\|^{2}]\leq\frac{(1-\bar{\alpha}_{T})T}{{\bar{\alpha}_{T}}}\delta$, where $\frac{(1-\bar{\alpha}_{T})T}{{\bar{\alpha}_{T}}}$ is the constant.
@@
This inequality holds when we use the DDIM reverse sampler and set the reverse step $s=1$. According to Song et al. (2020), the reverse process of DDIM is
$\hat{{\mathbf{x}}}_{\tau_{i-1}}=\sqrt{\bar{\alpha}_{\tau_{i-1}}}\left(\frac{\hat{{\mathbf{x}}}_{\tau_{i}}-\sqrt{1-\bar{\alpha}_{\tau_{i}}}{\bm{\epsilon}}_{\theta}(\hat{{\mathbf{x}}}_{\tau_{i}},\tau_{i})}{\sqrt{\bar{\alpha}_{\tau_{i}}}}\right)+\sqrt{1-\bar{\alpha}_{\tau_{i-1}}}{\bm{\epsilon}}_{\theta}(\hat{{\mathbf{x}}}_{\tau_{i}},\tau_{i}).$ (23)
In the discrete case (Ho et al., 2020), $\{\tau_{0},\ldots,\tau_{s}\}$ is a linearly increasing sub-sequence of $\{0,\ldots,T\}$, $\tau_{0}=0,\tau_{s}=T$. In the continuous case (Song et al., 2021), $\{\tau_{0},\ldots,\tau_{s}\}$ is a linearly increasing sequence in $[0,T]\subset[0.0,1.0]$, $\tau_{0}=0,0\leq\tau_{s}=T\leq 1$. $T$ is determined in the forward process. Setting the number of reverse steps $s=1$, then $\tau_{s}=T,\tau_{s-1}=\tau_{0}=0$, and it yields:
$\hat{{\mathbf{x}}}_{\tau_{s-1}}=\sqrt{\bar{\alpha}_{\tau_{s-1}}}\left(\frac{\hat{{\mathbf{x}}}_{\tau_{s}}-\sqrt{1-\bar{\alpha}_{\tau_{s}}}{\bm{\epsilon}}_{\theta}(\hat{{\mathbf{x}}}_{\tau_{s}},\tau_{s})}{\sqrt{\bar{\alpha}_{\tau_{s}}}}\right)+\sqrt{1-\bar{\alpha}_{\tau_{s-1}}}{\bm{\epsilon}}_{\theta}(\hat{{\mathbf{x}}}_{\tau_{s}},\tau_{s}),$ (24)
since $\hat{{\mathbf{x}}}_{\tau_{s}}=\hat{{\mathbf{x}}}_{T}={\mathbf{x}}_{T}^{a}$, we have
$\hat{{\mathbf{x}}}_{0}=\sqrt{\bar{\alpha}_{0}}\left(\frac{{\mathbf{x}}_{T}^{a}-\sqrt{1-\bar{\alpha}_{T}}{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{a},T)}{\sqrt{\bar{\alpha}_{T}}}\right)+\sqrt{1-\bar{\alpha}_{0}}{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{a},T),$ (25)
where since we use one-step reverse process, we reuse the notation of the final point of reverse process $\hat{{\mathbf{x}}}_{0}$ to represent the prediction of ${\mathbf{x}}_{0}$ from ${\mathbf{x}}_{T}^{a}$. Since $\sqrt{\bar{\alpha}_{0}}=1$, ${\mathbf{x}}_{T}^{a}=\sqrt{\bar{\alpha}_{T}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{T}}{\bm{\epsilon}}$, ${\mathbf{x}}_{0}^{a}={\mathbf{x}}_{0}+{\bm{\epsilon}}_{a}$, Eq. 25 can be written as
$\hat{{\mathbf{x}}}_{0}$ $={\mathbf{x}}_{0}^{a}+\frac{\sqrt{1-\bar{\alpha}_{T}}}{\sqrt{\bar{\alpha}_{T}}}({\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{a},T))$ (26)
$={\mathbf{x}}_{0}+\frac{\sqrt{1-\bar{\alpha}_{T}}}{\sqrt{\bar{\alpha}_{T}}}(\frac{\sqrt{\bar{\alpha}_{T}}}{\sqrt{1-\bar{\alpha}_{T}}}{\bm{\epsilon}}_{a}+{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{a},T)).$
@@
$\|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\|$ $=\frac{\sqrt{1-\bar{\alpha}_{T}}}{\sqrt{\bar{\alpha}_{T}}}\left\|\frac{\sqrt{\bar{\alpha}_{T}}}{\sqrt{1-\bar{\alpha}_{T}}}{\bm{\epsilon}}_{a}+{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{a},T)\right\|$ (27)
$=\frac{\sqrt{1-\bar{\alpha}_{T}}}{\sqrt{\bar{\alpha}_{T}}}\left\|\frac{\sqrt{\bar{\alpha}_{T}}}{\sqrt{1-\bar{\alpha}_{T}}}{\bm{\epsilon}}_{a}+{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{d},T)\right\|,$
where the second equivalence holds due to $t=T$ and ${\mathbf{x}}_{T}^{d}={\mathbf{x}}_{T}^{a}-k_{T}{\bm{\epsilon}}_{a}={\mathbf{x}}_{T}^{a}$. Considering that
$L_{b}$ $=\mathbb{E}_{{\bm{\epsilon}},t}\left[\left\|\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\cdot\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{a}+{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t)\right\|^{2}\right]$ (28)
$=\frac{1}{T}\sum_{t=0}^{T}\mathbb{E}_{{\bm{\epsilon}}}\left[\left\|\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\cdot\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{a}+{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{t}^{d},t)\right\|^{2}\right]$
@@
$\mathbb{E}_{{\bm{\epsilon}}}\left[\frac{{1-\bar{\alpha}_{T}}}{{\bar{\alpha}_{T}}}\left\|\frac{\sqrt{\bar{\alpha}_{T}}}{\sqrt{1-\bar{\alpha}_{T}}}{\bm{\epsilon}}_{a}+{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{d},T)\right\|^{2}\right]\leq T\cdot\delta$ (29)
@@
$\mathbb{E}_{{\bm{\epsilon}}}\left[\left|\hat{{\mathbf{x}}}_{0}-{\mathbf{x}}_{0}\right\|^{2}\right]$ $=\mathbb{E}_{{\bm{\epsilon}}}\left[\frac{{1-\bar{\alpha}_{T}}}{{\bar{\alpha}_{T}}}\left\|\frac{\sqrt{\bar{\alpha}_{T}}}{\sqrt{1-\bar{\alpha}_{T}}}{\bm{\epsilon}}_{a}+{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{T}^{d},T)\right\|^{2}\right]$ (30)
@@ C.3 The Proof of Theorem 2
The concept of infinite timestep can be viewed as dividing a finite length of time into infinitesimal intervals, which corresponds to the situation of the following SDEs proposed by Song et al. (2021). Denoting ${\mathbf{w}}$ as the standard Wiener process, $\bar{{\mathbf{w}}}$ as the reverse-time standard Wiener process, and $p_{t}({\mathbf{x}})$ the probability density of ${\mathbf{x}}_{t}$, the forward process can be described by
@@
$\propto\int\mathbbm{1}_{\{{\mathbf{x}}_{0}\notin{\mathbb{D}}_{a}\}}\exp\left(-\frac{\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\bm{\epsilon}}_{a}\|^{2}}{2(1-\bar{\alpha}_{t})}\right)p({\mathbf{x}}_{0}){\mathrm{d}}{\mathbf{x}}_{0}$
@@
$\left[\exp\left(-\frac{\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\bm{\epsilon}}_{a}\|_{2}^{2}}{2(1-\bar{\alpha}_{t})}\right)-\exp\left(-\frac{\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}\|_{2}^{2}}{2(1-\bar{\alpha}_{t})}\right)\right]{\mathrm{d}}{\mathbf{x}}_{0}.$
Therefore, we only need to compare $\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\bm{\epsilon}}_{a}\|_{2}^{2}$ and $\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}\|_{2}^{2}$ for comparing $P(B)$ and $P(D)$.
Since ${\mathbf{x}}_{t}^{a}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$ and ${\mathbf{x}}_{0}^{a}={\mathbf{x}}_{0}+{\bm{\epsilon}}_{a}$,
$\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\bm{\epsilon}}_{a}\|$ $=\|\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\bm{\epsilon}}_{a}\|$ (39)
$=\|\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}+k_{t}{\bm{\epsilon}}_{a}\|,$
$\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}\|$ $=\|\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}\|$ (40)
$=\|\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}+\sqrt{\bar{\alpha}_{t}}{\bm{\epsilon}}_{a}\|.$
@@
$\mathbb{E}_{{\bm{\epsilon}}}\|\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}+k_{t}{\bm{\epsilon}}_{a}\|<\mathbb{E}_{{\bm{\epsilon}}}\|\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}+\sqrt{\bar{\alpha}_{t}}{\bm{\epsilon}}_{a}\|.$ (41)
@@
$\mathbb{E}_{{\bm{\epsilon}}}\left[\exp\left(-\frac{\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}-(\sqrt{\bar{\alpha}_{t}}-k_{t}){\bm{\epsilon}}_{a}\|_{2}^{2}}{2(1-\bar{\alpha}_{t})}\right)\right]>\mathbb{E}_{{\bm{\epsilon}}}\left[\exp\left(-\frac{\|{\mathbf{x}}_{t}^{a}-\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}\|_{2}^{2}}{2(1-\bar{\alpha}_{t})}\right)\right],$ (42)
@@
Appendix D Additional Settings
D.1 Additional Training Settings of ADBM
We implemented ADBM based loosely on the original implementations by Song et al. (2021). Note that when fine-tuning the ADBM with the pre-trained checkpoint, following Song et al. (2021), we also adopted the continuous version of Eq. 9. This version is conceptually similar to its discrete counterpart, with the exception that $t$ in Eq. 9 represents a continuous value rather than a discrete one. Moreover, we used the Adam optimizer (Kingma & Ba, 2015) and incorporated the exponential moving average of models, with the average rate being 0.999. The batch size was set to 128 for SVHN and CIFAR-10, 112 for Tiny-ImageNet, and 64 for ImageNet-100 (due to memory constraints). All experiments were run using PyTorch 1.12.1 and CUDA 11.3 on 4 NVIDIA 3090 GPUs.
D.2 Additional Adaptive Attack Configurations
When performing the adaptive attacks (PGD with EOT) for the AP methods, we set the step sizes to be 0.007, 0.5, and 0.005 for $l_{\infty}$, $l_{1}$, and $l_{2}$ attacks, respectively. For the $l_{1}$ attack, we set the sparsity level as 0.95. As shown in Sec. 3, the adaptive evaluating is quite time-consuming. Thus following Nie et al. (2022), we conducted the adaptive attacks three times on a subset of 512 randomly sampled images from the test dataset and reported the mean accuracy along with the standard deviation. We note that sometimes the standard deviation is omitted to present the results more clearly.